OpenStudy (anonymous):

Find all polar coordinates of point P where P = (5, -pi/6) (5, -pi/6 + 2nπ) or (-5, -pi/6 + 2nπ) (5, -pi/6 + 2nπ) or (-5, -pi/6 + (2n + 1)π) (5, -pi/6 + 2nπ) or (5, pi/6 + (2n + 1)π) (5, -pi/6 + (2n + 1)π) or (-5, -pi/6+ 2nπ)

OpenStudy (anonymous):

@welshfella @Astrophysics

OpenStudy (amoodarya):

 hint : \[-\frac{\pi}{6}=-\frac{\pi}{6}+2kpi\]